Pronghorn Company purchased a delivery truck for $26,000 on January 1, 2020. The truck has an expected salvage value of $1,280, and is expected to be driven 103,000 miles over its estimated useful life of 10 years. Actual miles driven were 12,700 in 2020 and 10,200 in 2021.

a. Calculate depreciation expense per mile under units-of-activity method. (Round answer to 2 decimal places, e.g. 0.50.)

Depreciation expense $ per mile

b. Compute depreciation expense for 2020 and 2021 using (1) the straight-line method, (2) the units-of-activity method, and (3) the double-declining-balance method. (Round depreciation cost per unit to 2 decimal places, e.g. 0.50 and depreciation rate to 0 decimal places, e.g. 15%. Round final answers to 0 decimal places, e.g. 2,125.)

Depreciation Expense

2020

2021

(1) Straight-line method $ $
(2) Units-of-activity method $ $
(3) Declining-balance method $ $

c. Assume that Pronghorn uses the straight-line method. Prepare the journal entry to record 2020 depreciation.

d.Assume that Pronghorn uses the straight-line method. Show how the truck would be reported in the December 31, 2020, balance sheet. (Round answers to 0 decimal places, e.g. 2,125.)
Answer #1

(a) Depreciation expense per mile = (Cost - salvage value)/total miles

= (26,000 - 1,280)/103,000

= $0.24 per mile

(b) depreciation expense :-

2020 2021
Straight line method $2,472 $2,472
Units of activity method $3,048 $2,448
Double declining balance method $5,200 $4,160

Straight line method :-

Depreciation per year = (26,000 - 1,280)/10 = 2,472

Units of activity method:-

Depreciation expense 2020 = 12,700 x 0.24 = 3,048

Depreciation expense 2021 = 10,200 x 0.24 = 2,448

Double declining balance method :-

Depreciation rate = (1/10) x 2 = 20%

Depreciation expense 2020 = 26,000 x 20% = 5,200

Depreciation expense 2021 = (26,000 - 5,200) x 20% = 4,160

(c) Journal entry:-

Account Title and Explanation Debit Credit
Depreciation expense $2,472
Accumulated depreciation $2,472

(d) Balance sheet extract :-

Cost of Truck $26,000
Less: Accumulated depreciation ($2,472)
Book value $23,528